OpenStudy (yrivers36):

Solve the equation: 1(-x)+2x=3x+1. With explanation please.

OpenStudy (anonymous):

is that \[-x+2x=3x+1\]?

OpenStudy (yrivers36):

Sorry its 1/2(1-x)+2x=3x+1

OpenStudy (anonymous):

if so , comine like terms first on the left \[x=3x+1\] then subtract x from both sides \[0=2x+1\] subtract 1 from both sides \[-1=2x\] divide both sides by 2 \[-\frac{1}{2}=x\]

OpenStudy (callisto):

satellite73, how do you type fraction?

OpenStudy (yrivers36):

The answer should be -1/3